OpenStudy (anonymous):

A root of function f(x) is -1-2i. Which could be an equation for this function f(x)=x^2-2x-3 f(x)=x^2-2x+5 f(x)=x^2+2x-3 f(x)=X^2+2x+5

OpenStudy (sidsiddhartha):

x^2+2x+5

OpenStudy (anonymous):

@sidsiddhartha can you please tell me how you got that?

OpenStudy (whpalmer4):

you have a polynomial with only real numbers for its coefficients. That means that any complex roots come in pairs of the form \(a\pm bi\). You know one root is \(-1-2i\) so another root must be \(-1+2i\). Multiply \[(x-(-1-2i))(x-(-1+2i))\]and you've got your equation.

OpenStudy (whpalmer4):

That's a bit ugly, so there's a shortcut you can do: Say your roots are \(a\pm bi\), so the factors are \((x - a -bi)\) and \((x-a+bi)\): Instead of multiplying \[(x-a-bi)(x-a+bi)\]regroup it a bit as \[((x-a)-bi)((x-a)+bi)\]now let \(u = x-a\) and \(v=bi\) and you have\[(u-v)(u+v)\]which is a difference of squares: \[u^2-v^2 = (x-a)^2 - b^2i^2 = (x-a)^2 -b^2(-1) = (x-a)^2 + b^2\]You can expand that to \[(x^2-2ax+a^2) + b^2 = x^2+a^2+b^2-2ax\]which is easily evaluated: \[x^2 -2(-1)x +(-1)^2 + (-2)^2 = x^2+2x+1+4 = x^2+2x+5\]

OpenStudy (anonymous):

Use the discriminant. With the solution given, you need b^2-4ac to equal -16. That will give plus or minus 4i divided by two as the imaginary terms of the solutions. That means you are down to the second or fourth choices. From there, -b/2a has to be negative, so the fourth option is correct.

OpenStudy (anonymous):

@whpalmer4 I'm sorry, this is the first time I've seen a problem like this and I really don't know what you just did...thanks though

OpenStudy (whpalmer4):

That's fine, we can go back over it in smaller steps. Do you have a specific question you want me to answer, or should I just go through it with a different example that is a bit simpler?

OpenStudy (anonymous):

Yes please...where did the u's and v's come from and the bi's??

OpenStudy (whpalmer4):

Okay, the u's and v's were variables I introduced just to make it easier to see that you had a difference of squares. A difference of squares is when you have something like \[a^2-b^2\]but the letters are unimportant — it's <one thing>^2 - <another thing>^2 So I used u to represent (x-a), and v to represent bi Does that make sense?

OpenStudy (anonymous):

Yes it's starting to make sense, so like is a + over - bi the formula or something that will help with other problems

OpenStudy (whpalmer4):

okay, the \(a\pm bi\) is just shorthand for \(a+bi, a-bi\) Any complex number can be written in that form. \(i = \sqrt{-1}\)

OpenStudy (whpalmer4):

a number that has \(a = 0\) is called an imaginary number. a number that has \(b = 0\) is called a real number. a complex number is a combination of the two, so both \(a\) and \(b\) are not \(0\).

OpenStudy (anonymous):

So does it get any harder than this, like I know you have to add, subtract, and simplify and graph....I'm studying for my MCA's and I haven't learned this and it's going to be on the test but I think I'm starting to pick something up

OpenStudy (whpalmer4):

I don't know what MCA's are...

OpenStudy (anonymous):

Oh they stand for Minnesota Comprehension Assessment...it's our states test, it's the second most important test next to SAT and ACT. I'm taking the math test. But yeah thanks for all your help, I may not be a total expert yet but I think I understand it a little better.

OpenStudy (whpalmer4):

Ah, I'm from Minnesota too! But I'm living in sunny CA now. How did you enjoy this year's winter? I understand it may be over soon ;-)

OpenStudy (anonymous):

California! :o sooo jealous! this year winter was painful :'( thanks for reminding...But spring is coming slowing but surely D: lols but good for you for getting outta here while you can.

OpenStudy (whpalmer4):

the MCA postdates me...I left in the fall of 1981

OpenStudy (anonymous):

even more lucky...thanks again for your help

OpenStudy (whpalmer4):

do you have any more questions about this problem? or related problems?

OpenStudy (anonymous):

No that's the only one on our practice packet....but I can't figure out this one The surface area of a baseball is 177cm^2. What is the diameter of the baseball Use 3.14 for pi 3.75cm 7cm 7.5cm 14.1cm

OpenStudy (whpalmer4):

okay. do you know the formula for surface area of a sphere?

OpenStudy (anonymous):

A= 4 pi r 2.

OpenStudy (whpalmer4):

\[A = 4\pi r^2\] We know \(A=177 \text{ cm}^2\) Can you solve \[A = 4\pi r^2\] for \(r\) in terms of \(A\)?

OpenStudy (anonymous):

I got 12.82 177=4pier^2 177=12.26 r^2 -12.56 164.44= r^2 square root 164.44 on both sides 12.82

OpenStudy (whpalmer4):

No, solve the equation, then plug in numbers... \[A = 4\pi r^2\]divide both sides by 4\[\frac{A}{4} = \pi r^2\]Divide both sides by \(\pi\) \[\frac{A}{4\pi} = r^2\]Take square root of both sides \[r = \sqrt{\frac{A}{4\pi}}\]Now plug in numbers...

OpenStudy (whpalmer4):

by the way, it looks like you did this: \[A = 4\pi r^2\]\[177 = 4(3.14) r^2\]\[177 = 12.56r^2\] Then you subtracted 177-12.56...can't do that! \[177-12.56 = r^2\]

OpenStudy (anonymous):

I got it to solve for the diameter you divide the surface area for pi....and I got 7.5 cm

OpenStudy (whpalmer4):

You have to divide: \[177 = 12.56r^2\]\[\frac{177}{12.56} = \frac{12.56r^2}{12.56} = r^2\]

OpenStudy (whpalmer4):

but much better to do the problem with letters, then plug in numbers